OpenStudy (anonymous):

How much work does it take to move a 50 μC charge against a 12 V potential difference?

OpenStudy (anonymous):

work =V*Q =12*50*10^-6

OpenStudy (mayankdevnani):

the potential difference and the work per unit charge, done by an external agent are equal in magnitude so, W=(50 μC)(12V)=600μJ.

OpenStudy (mayankdevnani):

@franci2005